# Po-Heng Chou1, Chiapin Wang2, Kuan-Hao Chen2, and Wei-Chen Hsiao2

### v2

@@ IV-C Training Procedure
where $\delta_{t}=y_{t}-Q_{\theta}(\bm{s}_{t},\bm{w}_{t})$. Gradients are clipped with $\|\nabla_{\theta}\|\leq 5$ to prevent divergence, and $\epsilon$ decays exponentially from 1.0 to 0.01 during training. Convergence is typically achieved within 800 to 1000 episodes, after which the mean positioning error stabilizes below 0.5 m. Furthermore, the learned beam weights exhibit strong interpretability: high-SINR beams receive larger weights, while redundant or correlated beams are automatically suppressed, indicating that the agent implicitly learns the beam-geometry reliability without explicit CSI modeling.
@@ 13      Update $\theta\!\leftarrow\!\theta-\eta\nabla_{\theta}L_{t}(\theta)$ via Huber loss (17);
Algorithm 1 DQN-WLS Based Beam Positioning
@@ V-A Experimental Setup
All simulations were conducted in Python using the PyTorch 2.5 framework on a workstation equipped with an Intel i9 CPU and an NVIDIA RTX A6000 GPU. A 2-D coverage area of $1000\times 1000$ m2 is considered with $M=10$ visible beams. The discount factor is set to $\gamma=0.99$, the learning rate $\eta=10^{-3}$ (Adam optimizer), and the target network update interval $C=200$. Each training episode consists of $T=100$ steps, and the replay buffer stores $10^{4}$ transitions. Six algorithms are compared for performance evaluation:
  • PPO [14]: Proximal Policy Optimization baseline with a clipped surrogate objective.
  • DDQN [15]: Double Deep Q-Network that employs separate target and online networks to mitigate overestimation.
  • DQN [11]: Standard Deep Q-Network with experience replay and a single target estimator.
  • ALG-B [16]: Geometry-based intersection and exclusion algorithm for massive-MIMO LEO constellations, utilizing beam footprints and satellite identifiers to determine user position.
  • LSTM-WLS [17]: LSTM neural-network-based satellite weighting framework that predicts measurement reliability and applies adaptive WLS for single-epoch localization.
  • Proposed DQN-WLS: The proposed method integrates adaptive beam weighting with the WLS estimator to enable without explicit CSI estimation, real-time positioning.
The implementation of all six algorithms, including network configurations and simulation parameters, is available for reproducibility. [^1] Source code and configuration files are available at: https://github.com/BoneZhou/DRL-LEO-Beam-Positioning
@@ V-B Convergence Behavior
Fig. 1 illustrates the training convergence of cumulative reward and positioning error for the six algorithms across 1000 episodes. The PPO and DDQN models exhibit relatively unstable reward trajectories due to policy oscillation and overestimation bias, respectively. In particular, the DDQN agent shows limited improvement throughout training, which can be attributed to the quasi-continuous beam-weighting action space and the smooth reward landscape. Under such conditions, the value decoupling mechanism of DDQN provides limited benefit and may even slow convergence due to stale target updates, a phenomenon also observed in continuous control tasks. In contrast, the standard DQN achieves faster initial convergence but suffers from higher variance in later stages, indicating sensitivity to replay sampling. The proposed DQN-WLS demonstrates the most stable and rapid convergence, reaching a steady-state mean positioning error below 0.5 m within 800 episodes. This improvement stems from the adaptive beam-weighting mechanism that enables the agent to prioritize high-SINR beams while suppressing noisy or correlated ones. The geometry-guided regularization imposed by the WLS estimator further enhances training stability under dynamic multi-beam environments.
As a reference, the geometry-based ALG-B [16] and the LSTM-WLS [17] methods are not trained through episodic reinforcement learning. Hence, they do not appear in the convergence curves but are evaluated in subsequent subsections for accuracy and computational comparison. As illustrated in Fig. 1(d), the proposed agent achieves the highest reward stability and lowest error variance.
@@ V-C Positioning Visualization
Fig. 2 illustrates the final positioning results at Episode 1000 for three representative methods: the geometry-based ALG-B [16], the learning-based DQN [11], and the proposed DQN-WLS framework. As shown in Fig. 2(a), the ALG-B method produces a large localization deviation exceeding $10^{4}$ m, since it relies solely on fixed beam-intersection geometry without adaptive weighting or interference mitigation. The standard DQN in Fig. 2(b) reduces the error to the tens-of-meter level by learning partial beam reliability, but its positioning remains unstable due to the absence of explicit geometric constraints. In contrast, the proposed DQN-WLS shown in Fig. 2(c) achieves sub-meter accuracy (approximately $0.395$ m) with near-perfect alignment between the estimated and true UT positions. The visual comparison highlights that integrating reinforcement learning with the WLS estimator enables robust and high-precision positioning without explicit CSI estimation.
@@
Figure 2: Positioning visualization at Episode 1000 under identical beam geometry and noise conditions. The yellow star denotes the estimated UT position, and the black circle indicates the ground truth. Compared with the geometry-driven ALG-B and standard DQN, the proposed DQN-WLS aligns closely with the ground truth and yields the smallest residual error.
@@ V-D Performance Comparison
Table I summarizes the overall computational and localization performance of six representative algorithms, including both geometric and learning-based approaches. The conventional beam-intersection method (ALG-B [16]) exhibits the largest localization deviation (>3>3 km) due to its reliance on fixed beam footprints without adaptive weighting. Among DRL-based schemes, PPO [14] and DDQN [15] show unstable convergence and excessive computational overhead, while the standard DQN [11] significantly improves efficiency and reduces positioning error to 10.38 m. The LSTM-WLS [17] framework achieves extremely high accuracy (0.01 m) but incurs high training cost and longer inference latency, making it impractical for real-time onboard operation. By contrast, the proposed DQN-WLS achieves sub-meter accuracy (0.395 m) with the lowest runtime (9.43 s per episode) and nearly identical computational complexity to DQN ($\approx$19 GFLOPs). These results confirm that integrating adaptive beam weighting with WLS estimation enables superior accuracy–efficiency trade-offs suitable for resource-constrained LEO satellite payloads.
@@
Note: FLOPs (M) denote per-inference forward FLOPs. Total (G) includes the complete training forward and backward passes under identical episode and batch settings. RMSE values are averaged over 10 seeds with fixed geometries and interference profiles.
@@ V-E Discussion
The superior performance of DQN-WLS arises from its hybrid nature, combining the adaptivity of deep reinforcement learning with the interpretability and stability of model-based estimation. Through interaction with the environment, the agent learns to emphasize high-SINR and geometrically diverse beams while suppressing redundant or noisy ones. This implicit weighting behavior, realized without explicit CSI, ensures geometric consistency and robust convergence across varying satellite–user topologies. Compared with fully supervised models such as LSTM-WLS, DQN-WLS attains comparable accuracy but with over 80% lower computational cost, demonstrating its practicality for real-time onboard deployment in LEO constellations. The lightweight inference pipeline and small network footprint make DQN-WLS feasible for future satellite onboard processors.
@@ VI Conclusion
This paper proposed a DQN-WLS beam positioning framework for LEO satellite constellations that jointly optimizes beam weighting and user localization without explicit CSI estimation. Unlike geometric intersection or purely data-driven approaches, the proposed hybrid method leverages reinforcement learning for adaptive beam selection and integrates WLS for physics-consistent position estimation. Comprehensive simulations show that DQN-WLS reduces the mean positioning error by 99.3% compared with existing DRL baselines and achieves a near-real-time inference latency of less than 10 s. Given its low complexity, scalability, and interpretability, the proposed framework offers a promising foundation for future onboard and autonomous beam management in next-generation non-terrestrial networks.

### v3

@@ IV-C Training Procedure
where $\delta_{t}=y_{t}-Q_{\theta}(\bm{s}_{t},\bm{w}_{t})$. Gradients are clipped with $\|\nabla_{\theta}\|\leq 5$ to prevent divergence, and $\epsilon$ decays exponentially from 1.0 to 0.01 during training. Convergence is typically achieved within 800 to 1000 episodes in the considered settings, after which the positioning error enters a stable sub-meter regime. Furthermore, the learned satellite weights exhibit interpretable behavior: high-SINR beams tend to receive larger weights, while redundant or correlated beams are automatically suppressed, indicating that the agent learns beam reliability from pilot-level observations without explicit CSI modeling.
@@ 13      Update $\theta\!\leftarrow\!\theta-\eta\nabla_{\theta}L_{t}(\theta)$ via Huber loss (16);
Algorithm 1 DQN-WLS Based Satellite Positioning
@@ V-A Experimental Setup
All simulations were conducted in Python using the PyTorch 2.5 framework on a workstation equipped with an Intel i9 CPU and an NVIDIA RTX A6000 GPU. A 2-D coverage area of $1000\times 1000$ m2 is considered with $M=10$ visible satellites. The discount factor is set to $\gamma=0.99$, the learning rate $\eta=10^{-3}$ (Adam optimizer), and the target network update interval $C=200$. Each training episode consists of $T=100$ steps, and the replay buffer stores $10^{4}$ transitions.
Residual Doppler. After ephemeris-based Doppler pre-compensation, the residual Doppler is modeled as a zero-mean Gaussian random variable with variance $\sigma_{\Delta f_{\mathrm{D}}}^{2}=5$ (Hz2), and its impact is incorporated into the measurement reliability via (7)–(8), where $T_{p}=N_{p}T_{s}$.
We compare six methods: PPO [14], DDQN [15], DQN [11], ALG-B [16], LSTM-WLS [17], and the proposed DQN-WLS. The implementation of all six algorithms, including network configurations and simulation parameters, is available for reproducibility. [^1] Source code and configuration files are available at: github.com/BoneZhou/DRL-LEO-Beam-Positioning
@@ V-B Convergence Behavior
Fig. 2 reports the learning curves of PPO, DDQN, DQN, and the proposed DQN-WLS. The proposed DQN-WLS converges stably and achieves the lowest positioning error among the evaluated DRL baselines, benefiting from the geometry-consistent WLS refinement.
@@ V-C Positioning Visualization
Fig. 3 compares ALG-B, DQN, and the proposed DQN-WLS at Episode 1000. The proposed DQN-WLS closely matches the ground truth and achieves sub-meter accuracy (0.395 m RMSE), while ALG-B and DQN exhibit much larger deviations.
@@
Figure 3: Positioning visualization at Episode 1000 under identical satellite geometry and noise conditions. The yellow star denotes the estimated UT position, and the black circle indicates the ground truth. Compared with the geometry-driven ALG-B and standard DQN, the proposed DQN-WLS aligns closely with the ground truth and yields the smallest residual error.
@@ V-D Performance Comparison
Table I summarizes the overall computational and localization performance of six representative algorithms, including both geometric and learning-based approaches. The conventional beam-intersection method (ALG-B [16]) exhibits the largest localization deviation (>3>3 km) due to its reliance on fixed beam footprints without adaptive weighting. Among DRL-based schemes, PPO [14] and DDQN [15] incur substantially higher runtime under the considered training configuration, while the standard DQN [11] improves efficiency and reduces the positioning error to 10.38 m. The LSTM-WLS [17] framework achieves extremely high accuracy (0.01 m) but incurs high training cost and longer inference latency, making it impractical for real-time onboard operation. By contrast, the proposed DQN-WLS achieves sub-meter accuracy (0.395 m) with the lowest runtime (9.43 s per episode) and nearly identical computational complexity to DQN ($\approx$19 GFLOPs). These results confirm that the proposed DQN-WLS achieves a favorable accuracy–efficiency trade-off with low per-step complexity. In addition, we include an ablation where ${\Delta f_{\mathrm{D}}}_{i}=0$ (Doppler-free) and observe only a negligible RMSE change in our considered setting, indicating that the proposed method is robust under typical residual Doppler after compensation.
@@
Note: Time and RMSE are averaged over 10 seeds under identical settings.
@@ V-E Discussion
The favorable accuracy-runtime trade-off of DQN-WLS stems from its hybrid structure, combining the adaptivity of reinforcement learning with the stability and interpretability of model-based estimation. Through interaction with the environment, the agent learns to emphasize high-SINR and geometrically informative beams while suppressing noisy or redundant ones, and the augmented WLS solver enforces geometric consistency during localization. We also include supervised LSTM-based weighting (LSTM-WLS [17]) as a reference that targets accuracy-optimal estimation in its supervised setting. In contrast, the proposed DQN-WLS is designed for CSI-free and deployment-oriented operation with low runtime overhead, which is relevant to resource-constrained LEO payload scenarios.
@@ VI Conclusion
This paper investigated a DQN-WLS framework for CSI-free multi-satellite positioning in LEO constellations that learns discrete satellite-weighting policies from pilot-level observations and geometric features. By coupling a lightweight DQN policy with an augmented WLS estimator, the proposed method provides a deployment-oriented accuracy-runtime trade-off and preserves physics-consistent localization with clock-bias estimation. Simulation results in representative settings demonstrate sub-meter accuracy with low computational overhead, suggesting that hybrid RL-WLS designs are a promising direction for CSI-free beam-based positioning in NTNs. Extending the framework to broader beam counts, richer interference models, and more diverse geometries is a natural next step.